# Are They the Same? Exploring Visual Correspondence Shortcomings of Multimodal LLMs

### v2

@@ 5 Experiments
Table 1: MMVM Benchmark results. Accuracy is the metric, and the overall accuracy is computed across all 1,510 evaluation samples. The accuracy for each of the eight match types is calculated separately on their respective samples. The full term of the match type abbreviation can be found in Sec. 3.1. For MLLMs that only support single-image input, we simply concatenate all the images vertically into one image and then input it.
@@
4B CoLVA-InternVL2-4B (Ours) 49.87 43.21 41.38 45.79 77.92 44.43 53.60 44.89 53.24
@@
Baselines and Datasets. We use the pre-trained InternVL2 4B [12] as the baseline. Please refer to the appendix for more experimental results of CoLVA built on different baseline MLLMs. During the SFT phase, we utilize the LLaVA SFT data [49] and our MMVM dataset. The LLaVA SFT data comprises approximately 665k conversation entries, and our MMVM data includes around 220k.
Implementation Details. Our model comprises three components: a pre-trained MLLM InternVL2-4B [12], a fine-grained vision expert RADIO [64], and a RADIO adapter. We adopt Xtuner [13] codebase to implement our method. We maintain the original architecture of both InternVL2-4B and RADIO, while the RADIO adapter is implemented using a two-layer MLP. Our training includes two stages: pre-training and supervised fine-tuning (SFT). We freeze the MLLM and RADIO during the pre-train stage, focusing solely on training the RADIO Adapter. During the SFT stage, we freeze the RADIO, the RADIO adapter, and all components of InternVL2-4B except the LLM. The LLM of the MLLM is tuned applying LoRA [26]. We use the Osprey-727K [99] for the region-text pre-traning.
5.1 Main Results
Results on MMVM benchmark. As shown in Tab. 1, we report the average accuracy of over 20 open-source MLLMs of varying sizes, three proprietary MLLMs, and our method on the MMVM Benchmark. On the MMVM Benchmark, none of the open-source or proprietary MLLMs achieved an overall accuracy exceeding 50%. By introducing object-level contrastive learning, fine-grained vision expert, and instruction augmentation, our method achieved significant performance improvements, reaching state-of-the-art performance and surpassing the previous highest accuracy obtained by GPT4o. Among all open-source MLLMs, InternVL2 [12] achieved the highest accuracy in the sub-4B and 13Bsimilar-$\sim$40B tiers, while Qwen2-VL [81] excelled in the 4Bsimilar-$\sim$13B and above-40B tiers. Overall, Qwen2-VL-72B [81] achieved the highest accuracy among all open-source MLLMs, approaching the accuracy of the proprietary GPT4o (38.08 vs. 42.65).
5.2 Ablation Study and Analysis
The Effectiveness of Data and Methods. As shown in Tab. 2, we used InternVL2-4B as our base model, achieving an overall accuracy of 17.62% on our MMVM benchmark. By fine-tuning InternVL2-4B with LLaVA SFT data [49] and our MMVM data, we observed a significant increase in overall accuracy (+14.76%), validating the effectiveness of our MMVM SFT data. We adopted this fine-tuned InternVL2-4B as a strong baseline and integrated our methods, which include object-level contrastive learning, fine-grained vision expert, and instruction augmentation. By incorporating the fine-grained vision expert into the fine-tuned InternVL2 and using object-level contrastive learning to pre-train its adapter, we observed an 8.07% improvement in overall accuracy. We believe that VE provides the basic knowledge necessary for object matching, such as the recognition of appearance, position, size, and other attributes. OCL offers an appropriate training strategy to enable MLLM to comprehend the knowledge provided by VE. Sufficient object matching data is crucial for teaching the MLLM to utilize this knowledge to perform object matching. Due to the substantial gap between the feature space of the fine-grained vision expert and that of the MLLM, directly using the visual features from the fine-grained vision expert did not yield any significant impact (32.25 vs. 32.38). Notably, directly applying object-level contrastive learning to the visual encoder of InternVL2 resulted in only a limited improvement in overall accuracy (34.05 vs. 32.38), as the CLIP-style vision backbone lacks fine-grained visual features. Further augmentation of instructions led to an additional accuracy gain of 5.38%.
The Effectiveness of Object-level Contrastive Learning. Our method employs object-level contrastive learning to pre-train the RADIO adapter. Compared to other standard methods that use image-text pairs (Image-Text) or region-text pairs (Region-Text) to pre-train the adapter by applying autoregressive training objective, our method (Region-Region) demonstrates significant advantages (45.83 vs. 33.64, or 45.83 vs. 30.93).
The Alternatives of RADIO. Our method remains effective for vision-only self-supervised learning (SSL) models. In particular, we replaced RADIO with DINOv2 [59] and ConvNext-L [84]. As shown in Tab. 4, our method still proves effective for vision-only SSL models, with a significant improvement in accuracy (40.34 vs 32.38). However, there is a gap between RADIO and DINOv2. This means both a semantic and spatial-aware visual expert is needed to achieve better results. CoLVA with ConvNext CLIP demonstrates a superior understanding of text markers (TM) compared to DINOv2 and RADIO but exhibits worse overall performance.
Data OCL VE IA OA $\Delta$
17.62 -
32.38 +14.76
34.05 +1.67
32.25 -0.13
40.45 +8.07
45.83 +5.38
Table 3: The effectiveness of object-level contrastive loss. Image-Text/Region-Text means using image-text/region-text pairs for align pre-training. Region-Region means using contrastive loss for align pre-training.
Metric Baseline No Alignment Image-Text Region-Text Region-Region
Overall Acc. 32.38 32.25 33.64 30.93 40.45
Overall CL SP TM SZ RP OO BR OM
Baseline 32.38 25.04 24.14 32.71 74.03 19.00 35.20 43.18 36.57
RADIO 45.83 38.30 31.03 41.12 76.62 41.71 51.20 39.77 46.76
DINOv2 40.34 33.72 44.83 42.06 64.94 32.28 36.00 35.80 39.81
ConvNext-L 39.80 31.59 34.48 42.99 77.92 26.57 48.80 44.32 44.44
5.3 Generalization study of CoLVA
MLLM CoLVA OA CL SP TM SZ RP OO BR OM
InternVL2-4B $\times$ 17.62 14.73 34.48 17.76 15.58 10.28 24.00 31.25 21.30
45.83 38.30 31.03 41.12 76.62 41.71 51.20 39.77 46.76
Qwen2VL-2B $\times$ 15.69 13.42 20.69 17.75 31.16 9.57 22.40 18.75 16.67
47.48 40.92 31.03 47.66 68.83 50.57 49.60 33.52 38.42
LLaVA1.5-7B $\times$ 14.64 12.44 10.34 16.82 29.87 8.00 19.20 19.32 16.20
36.56 29.13 27.59 38.32 71.43 26.14 40.80 42.04 39.35
Table 6: The Impact of CoLVA on the Common VQA Capabilities of MLLMs.
MLLM CoLVA MMBench DEV MME MME MMStar MMMU Val POPE BLINK NaturalBench NaturalBench NaturalBench VideoRefer-Bench
Overall Perception Reasoning Overall Overall Overall Overall Q_Acc I_Acc G_Acc Average
InternVL2-4B $\times$ 77.40 1536.14 533.93 54.40 47.56 84.52 45.76 44.71 48.63 19.52 60.91
77.32 1552.82 549.64 53.47 44.11 86.11 47.24 47.89 52.16 20.84 62.94
Exploration CoLVA on more MLLMs. To validate the generalization capability of our CoLVA across different MLLMs, we integrated CoLVA into three distinct MLLMs: InternVL2-4B [12], Qwen2VL-2B [81], and LLaVA1.5-7B [49]. Their performance on the MMVM benchmark is presented in Tab. 5. The results demonstrate that CoLVA significantly enhances the fine-grained visual matching capabilities across all three MLLMs. Notably, LLaVA1.5-7B, which has not undergone multi-image training, exhibited the smallest accuracy improvement after integrating CoLVA. In contrast, both InternVL2-4B and Qwen2VL-2B, having been trained with multiple images, showed substantial accuracy improvements on the MMVM benchmark post-CoLVA integration. The most significant accuracy enhancement was observed with Qwen2VL-2B, likely due to its unique architectural innovation, the Multimodal Rotary Position Embedding (M-RoPE). This feature enables MLLMs to simultaneously capture and integrate positional information from one-dimensional text sequences, two-dimensional visual images, and three-dimensional videos, facilitating fine-grained visual matching tasks within spatiotemporally varying multi-image contexts.
Results on common VQA benchmarks. To investigate whether CoLVA adversely affects the inherent general visual question answering (VQA) capabilities of MLLMs, we conducted tests across six relevant benchmarks: MMBench [51], MME [18], MMStar [9], MMMU [101], POPE [39], BLINK [20], NaturalBench [33], and VideoRefer-Bench [100]. The results are presented in Tab. 6. We used InternVL2-4B as the baseline and integrated our CoLVA into this framework. The results indicate that the negative impact of CoLVA on the general VQA capabilities of MLLMs is minimal. In fact, it even shows positive effects on the MME, POPE, BLINK, NaturalBench, and VideoRefer-Bench benchmarks. Therefore, our CoLVA does not compromise the original general VQA capabilities of MLLMs and can be a good supplementary of current mainstream VQA datasets.
6 Conclusion
In this paper, we present the MMVM benchmark and corresponding evaluation metrics to comprehensively evaluate current MLLMs’ performance in visual correspondence. The experiment results demonstrate all MLLMs perform poorly, with none achieving accuracy above 50%, including GPT-4o. To address the significant deficiencies of current MLLMs in visual correspondence, we designed an automatic annotation pipeline to generate a 220K visual matching SFT dataset with reasoning. Furthermore, we developed CoLVA through two novel designs: combining object-level contrastive learning with RADIO to obtain descriptive and fine-grained visual features and an instruction augmentation strategy for better training. Experiments demonstrate that our novel designs bring a performance improvement of 13.45 OA. Benefiting from our SFT data and the novel designs, our proposed CoLVA-Qwen2VL-7B achieves 51.06 OA on the MMVM benchmark, which surpasses the baseline Qwen2-VL-7B-Instruct with a 23.58 OA performance improvement.
Data OCL VE IA OA CL SP TM SZ RP OO BR OM
17.62 14.73 34.48 17.76 15.58 10.28 24.00 31.25 21.30
32.38 25.04 24.14 32.71 74.03 19.00 35.20 43.18 36.57
34.05 25.78 26.77 31.97 75.01 22.32 35.29 42.98 37.51
32.25 24.22 27.59 31.78 68.83 19.14 35.20 40.34 39.35
40.45 33.72 44.85 39.37 75.33 30.00 48.00 38.65 44.78
45.83 38.30 31.03 41.12 76.62 41.71 51.20 39.77 46.76
Table 8: The impact of Qwen2VL-CoLVA on general benchmarks.
MLLM CoLVA MME MME POPE BLINK
perception reasoning Overall Overall
Qwen2VL-2B $\times$ 1471.10 404.64 86.83 44.50
1540.14 418.57 88.01 46.98
Appendix A More Experiment Result
Appendix B More information on MMVM Benchmark
The minimum, maximum, and average number of choices per sample are 2, 37, and 10, respectively. We perform the ablation study on the number of choices, and the results are presented in Tab. 10. Fewer options make it easier for the model to correctly match the target.
Table 9: The split of MMVM benchmark.
Method Total In-domain split Out-domain split
GPT4o 42.65 46.46 38.47
InternVL2-4B 17.62 21.01 13.89
CoLVA-4B 49.87 57.22 41.67
Table 10: Number of choices.
k Overall
4 59.14
10 51.32
all 49.87
Figure 8: Potential real-world applications of CoLVA
Appendix C Potential real-word applications of CoLVA
Appendix D More Implementation Details
More training details. Our CoLVA training is divided into two phases: the pre-training phase and the fine-tuning phase. During the pre-training phase, we sample 500k images with segmentation labels from SA1B [30]. For each image, we apply augmentations such as Crop, Resize, Flip, and Rotation to simulate a pseudo video. We then sample two frames from this pseudo video to serve as our training samples. Taking InternVL2 [12] as the base model and RADIO [64] as the vision expert, we input one image into the InternVL2 visual encoder and the other into RADIO. When selecting the (anchor, positive, negatives) triplet, the anchor is chosen from the image features output by RADIO, while the positive and negatives are selected from the image features output by the InternVL2 visual encoder. We perform full training from scratch on the RADIO adapter using only object-level contrastive loss.
Appendix E More visualization results
Appendix F Further Discussion

### v3

@@ 5 Experiments
Table 1: MMVM Benchmark Results. Given that CL and RP are the two primary matching cues, we report the overall accuracy, CL accuracy, and RP accuracy under four different settings: with 4, 8, 12, and all candidate options. The overall accuracy is computed across all 1,510 evaluation samples. CL accuracy and RP accuracy are calculated separately on their respective samples. The full terms of the matching type abbreviations can be found in Sec. 3.2. For MLLMs that only support single-image input, we simply concatenate all the images vertically into a single image and then provide it as input.
Model Size Method Overall CL RP Overall-12 CL-12 RP-12 Overall-8 CL-8 RP-8 Overall-4 CL-4 RP-4
$\sim$4B Qwen2-VL-2B-Instruct [83] 15.69 13.42 9.57 16.82 14.57 11.71 19.93 17.68 16.28 32.35 28.97 30.86
DeepSeek-VL-1.3B [54] 16.82 12.60 10.43 17.46 13.97 10.70 21.77 18.36 19.18 28.21 25.96 23.58
InternVL2-4B [12] 17.62 14.73 10.28 18.34 18.17 14.86 20.73 20.78 18.71 35.76 35.84 35.00
4Bsimilar-$\sim$13B DeepSeek-VL-7b [54] 17.68 14.24 10.00 19.22 14.79 11.98 23.04 18.70 16.29 27.01 24.13 20.56
LLaVA-Next-Interleave-7B [36] 19.34 15.88 10.71 21.03 17.75 12.86 25.01 19.53 16.64 28.47 26.04 20.36
LLaVA-OneVision-ov-7B [35] 20.92 16.69 14.28 22.01 17.96 15.47 25.85 21.30 19.67 29.87 25.46 23.33
Qwen2-VL-7B-Instruct [83] 27.48 24.87 17.85 28.34 25.51 18.97 30.04 26.98 22.11 36.75 33.21 27.30
13Bsimilar-$\sim$40B LLaVA-Next-34B [52] 15.03 11.29 8.71 15.86 12.33 9.22 19.79 16.84 11.79 24.97 21.46 17.83
VILA1.5-40B [46] 15.36 14.73 5.00 16.95 15.60 6.45 20.85 18.78 12.01 26.12 24.37 17.33
InternVL2-40B [12] 26.03 24.88 16.86 27.99 26.13 19.46 32.63 30.70 24.97 37.35 34.93 30.05
40Bsimilar-$\sim$ InternVL2-76B [12] 25.83 24.06 19.28 27.56 26.13 21.99 32.76 30.07 26.94 36.17 34.96 30.54
LLaVA-OneVision-ov-72B [35] 29.34 28.48 21.14 29.89 29.31 23.14 32.55 33.03 27.34 37.43 35.20 30.34
InternVL2.5-78B [11] 36.42 35.02 25.86 39.01 35.72 33.40 42.11 39.31 34.79 45.89 43.41 37.57
Qwen2-VL-72B-Instruct [83] 38.08 37.64 32.28 39.77 35.00 31.94 42.31 39.83 35.62 47.68 45.23 39.23
Unkown Claude3-5V-Sonnet 40.20 34.21 34.86 41.75 33.98 37.01 45.77 38.02 41.39 51.35 43.79 48.63
GeminiPro1-5 40.73 36.00 35.14 43.01 39.27 36.98 46.66 41.97 38.07 52.62 50.30 45.13
GPT4o-20240806 42.65 39.28 32.28 44.71 43.63 39.65 49.30 45.37 42.11 56.76 53.24 47.57
4B CoLVA-InternVL2-4B (Ours) 49.80 42.72 44.86 51.06 44.19 46.86 53.38 46.48 49.71 59.47 51.72 57.86
Table 2: The impact of CoLVA on the single-image VQA capabilities of MLLMs.
MLLM CoLVA MMBench DEV MME MME MMStar MMMU Val POPE BLINK
Overall Perception Reasoning Overall Overall Overall Overall
InternVL2-4B [12] $\times$ 77.40 1536.14 533.93 54.40 47.56 84.52 45.76
77.32 1552.82 549.64 53.47 44.11 86.11 47.24
Table 3: The impact of CoLVA on the multi-image VQA capabilities of MLLMs.
MLLM CoLVA NaturalBench NaturalBench NaturalBench VideoRefer-Bench
Q_Acc I_Acc G_Acc Average
InternVL2-4B [12] $\times$ 44.71 48.63 19.52 60.91
47.89 52.16 20.84 62.94
Baselines and Datasets. We use the pre-trained InternVL2 4B [12] as the baseline. During the SFT phase, we utilize the LLaVA SFT data [51] and our MMVM dataset. The LLaVA SFT data comprises approximately 665k conversation entries, and our MMVM data includes around 220k. Please note that all ablation experiments were conducted using 30% of these data.
Implementation Details. Our model comprises three components: a pre-trained MLLM InternVL2-4B [12], a fine-grained vision expert RADIO [66], and a RADIO adapter. We adopt Xtuner [13] codebase to implement our method. Please refer to the appendix for the details.
5.1 Main Results
Results on MMVM benchmark. As shown in Tab. 1, we report the average accuracy of multiple open-source MLLMs of varying sizes, three proprietary MLLMs, and our method on the MMVM Benchmark. In the MMVM benchmark, none of the open-source or proprietary MLLMs achieved an overall accuracy exceeding 50% under the setting of all choice options. Compared to relative position in the scene, these MLLMs demonstrate a stronger capability in perceiving and utilizing color as a matching cue, as evidenced by the CL accuracy consistently surpassing the RP accuracy across almost all settings. By introducing object-level contrastive learning, fine-grained vision expert, and instruction augmentation, our method achieved significant performance improvements, reaching state-of-the-art performance and surpassing the previous highest accuracy obtained by GPT4o. Among all open-source MLLMs, InternVL2 [12] achieved the highest accuracy in the sub-4B and 13Bsimilar-$\sim$40B tiers, while Qwen2-VL [83] excelled in the 4Bsimilar-$\sim$13B and above-40B tiers. Overall, Qwen2-VL-72B [83] achieved the highest accuracy among all open-source MLLMs, approaching the accuracy of the proprietary GPT4o (38.08 vs. 42.65).
Results on common VQA benchmarks. To investigate whether CoLVA adversely affects the inherent general visual question answering (VQA) capabilities of MLLMs, we conducted tests across six relevant benchmarks: MMBench [53], MME [19], MMStar [9], MMMU [103], POPE [41], BLINK [21], NaturalBench [34], and VideoRefer-Bench [102]. The results are presented in Tab. 2 and Tab. 3. We used InternVL2-4B as the baseline and integrated our CoLVA into this framework. The results indicate that the negative impact of CoLVA on the general VQA capabilities of MLLMs is minimal. In fact, it even shows positive effects on the MME, POPE, BLINK, NaturalBench, and VideoRefer-Bench benchmarks. Therefore, our CoLVA does not compromise the original general VQA capabilities of MLLMs and can be a good supplement to current mainstream VQA datasets.
5.2 Ablation Study and Analysis
The Effectiveness of Data and Methods. As shown in Tab. 4, we used InternVL2-4B as our base model, achieving an overall accuracy of 17.62% on our MMVM benchmark. By fine-tuning InternVL2-4B with LLaVA SFT data [51] and our MMVM data, we observed a significant increase in overall accuracy (+14.76%), validating the effectiveness of our MMVM SFT data. We adopted this fine-tuned InternVL2-4B as a strong baseline and integrated our methods, which include object-level contrastive learning, fine-grained vision expert, and instruction augmentation. By incorporating the fine-grained vision expert into the fine-tuned InternVL2 and using object-level contrastive learning to pre-train its adapter, we observed an 8.07% improvement in overall accuracy. We believe that VE provides the basic knowledge necessary for object matching, such as recognizing appearance, position, size, and other attributes. OCL offers an appropriate training strategy to enable MLLM to comprehend the knowledge provided by VE. Sufficient object matching data is crucial for teaching the MLLM to utilize this knowledge to perform object matching. Due to the substantial gap between the feature space of the fine-grained vision expert and that of the MLLM, directly using the visual features from the fine-grained vision expert did not yield any significant impact (32.25 vs. 32.38). Notably, directly applying object-level contrastive learning to the visual encoder of InternVL2 resulted in only a limited improvement in overall accuracy (34.05 vs. 32.38), as the CLIP-style vision backbone lacks fine-grained visual features. Further augmentation of instructions led to an additional accuracy gain of 5.38%.
The Effectiveness of Object-level Contrastive Learning. Our method employs object-level contrastive learning to pre-train the RADIO adapter. As shown in Tab. 5, compared to other standard methods that use image-text pairs (Image-Text) or region-text pairs (Region-Text) to pre-train the adapter by applying autoregressive training objective, our method (Region-Region) demonstrates significant advantages (40.45 vs. 33.64, or 40.45 vs. 30.93).
The Alternatives of RADIO. Our method still works well for vision self-supervised learning models. In particular, we replaced RADIO with DINOv2 [61] and ConvNext-L [86]. As shown in Tab. 6, our method still proves effective for vision-only SSL models, with a significant improvement in accuracy (40.34 vs 32.38). However, there is a gap between RADIO and DINOv2. This means both a semantic and spatial-aware visual expert is needed to achieve better results. CoLVA with ConvNext CLIP demonstrates a superior understanding of text markers (TM) compared to DINOv2 and RADIO but exhibits worse overall performance.
Data OCL VE IA OA $\Delta$
17.62 -
32.38 +14.76
34.05 +1.67
32.25 -0.13
40.45 +8.07
45.83 +5.38
Table 5: The effectiveness of object-level contrastive loss. Image-Text/Region-Text means using image-text/region-text pairs for pre-training. Region-Region means using contrastive loss for pre-training.
Metric Baseline No Alignment Image-Text Region-Text Region-Region
Overall Acc. 32.38 32.25 33.64 30.93 40.45
Overall CL SP TM SZ RP OO BR OM
Baseline 32.38 25.04 24.14 32.71 74.03 19.00 35.20 43.18 36.57
RADIO [66] 45.83 38.30 31.03 41.12 76.62 41.71 51.20 39.77 46.76
DINOv2 [61] 40.34 33.72 44.83 42.06 64.94 32.28 36.00 35.80 39.81
ConvNext-L [86] 39.80 31.59 34.48 42.99 77.92 26.57 48.80 44.32 44.44
5.3 Generalization study of CoLVA
MLLM CoLVA OA CL RP
InternVL2-4B [12] $\times$ 17.62 14.73 10.28
45.83 38.30 41.71
Qwen2VL-2B [83] $\times$ 15.69 13.42 9.57
47.48 40.92 50.57
LLaVA1.5-7B [49] $\times$ 14.64 12.44 8.00
36.56 29.13 26.14
To validate the generalization of CoLVA on different MLLMs, we integrate CoLVA into three distinct MLLMs: InternVL2-4B [12], Qwen2VL-2B [83], and LLaVA1.5-7B [51]. Their performance on the MMVM benchmark is presented in Tab. 7. The results demonstrate that CoLVA significantly improves the fine-grained visual matching capabilities across all three MLLMs. Notably, LLaVA1.5-7B, which has not undergone multi-image training, exhibited the smallest accuracy improvement after integrating CoLVA. In contrast, both InternVL2-4B and Qwen2VL-2B, having been trained with multiple images, showed substantial accuracy improvements on the MMVM benchmark with our CoLVA integration.
6 Conclusion
This paper presents the MMVM benchmark, the first corresponding fine-grained visual correspondence evaluation benchmark for current MLLMs. The results demonstrate all MLLMs perform poorly, with none achieving accuracy above 50% under the setting of all candidate options, including GPT-4o. To address the significant weakness of current MLLMs in visual correspondence, we design an automatic annotation pipeline to generate a 220K visual matching SFT dataset with reasoning. Furthermore, we propose CoLVA through two novel designs: combining object-level contrastive learning with RADIO to obtain descriptive visual features and an instruction augmentation strategy. Experiments demonstrate that our novel designs improve base MLLM by 13.45 OA. Benefiting from our SFT data and the novel designs, our proposed CoLVA-InternVL2-4B achieves 49.80 OA on the MMVM benchmark, surpassing the baseline InternVL2-4B with a 32.18 OA performance improvement.
Data OCL VE IA OA CL SP TM SZ RP OO BR OM
17.62 14.73 34.48 17.76 15.58 10.28 24.00 31.25 21.30
32.38 25.04 24.14 32.71 74.03 19.00 35.20 43.18 36.57
34.05 25.78 26.77 31.97 75.01 22.32 35.29 42.98 37.51
32.25 24.22 27.59 31.78 68.83 19.14 35.20 40.34 39.35
40.45 33.72 44.85 39.37 75.33 30.00 48.00 38.65 44.78
45.83 38.30 31.03 41.12 76.62 41.71 51.20 39.77 46.76
Table 9: The impact of Qwen2VL-CoLVA on general benchmarks.
MLLM CoLVA MME MME POPE BLINK
perception reasoning Overall Overall
Qwen2VL-2B $\times$ 1471.10 404.64 86.83 44.50
1540.14 418.57 88.01 46.98
Table 10: The split of MMVM benchmark.
Method Total In-domain split Out-domain split
GPT4o 42.65 46.46 38.47
InternVL2-4B 17.62 21.01 13.89
CoLVA-4B 49.87 57.22 41.67
Appendix A More Experiment Result
Figure 8: Potential real-world applications of CoLVA
Appendix B More information on MMVM Benchmark
Appendix C Potential real-world applications of CoLVA
Appendix D More Implementation Details
More training details. Our model comprises three components: a pre-trained MLLM InternVL2-4B [12], a fine-grained vision expert RADIO [66], and a RADIO adapter. We adopt Xtuner [13] codebase to implement our method. We maintain the original architecture of both InternVL2-4B and RADIO, while the RADIO adapter is implemented using a two-layer MLP. Our training includes two stages: pre-training and supervised fine-tuning (SFT). We freeze the MLLM and RADIO during the pre-train stage, focusing solely on training the RADIO Adapter. During the SFT stage, we freeze the RADIO, the RADIO adapter, and all components of InternVL2-4B except the LLM. The LLM of the MLLM is trained by applying LoRA [27].
During the pre-training phase, we sample 500k images with segmentation labels from SA1B [31]. For each image, we apply augmentations such as Crop, Resize, Flip, and Rotation to simulate a pseudo video. We then sample two frames from this pseudo video to serve as our training samples. Taking InternVL2 [12] as the base model and RADIO [66] as the vision expert, we input one image into the InternVL2 visual encoder and the other into RADIO. When selecting the (anchor, positive, negatives) triplet, the anchor is chosen from the image features output by RADIO, while the positive and negatives are selected from the image features output by the InternVL2 visual encoder. We perform full training from scratch on the RADIO adapter using only object-level contrastive loss.
Appendix E More visualization results
More PCA visualizations. In Fig. 9, we present additional PCA visualizations. The results reveal that the matched target (represented by a red dot) and other candidate objects (represented by blue dots) are clustered together, while being distant from the query object (represented by a red star). This clustering pattern makes it challenging for InternVL2 to distinguish the correct object. In contrast, our CoLVA brings the matched target and the query object closer together while distancing them from other candidate objects. This indicates that our CoLVA has learned fine-grained and discriminative visual features, which are beneficial for visual matching tasks.
Appendix F Further Discussion
Table 11: More MMVM Benchmark results. Accuracy is the metric, and the overall accuracy is computed across all 1,510 evaluation samples. The accuracy for each of the eight match types is calculated separately on their respective samples. The full term of the match type abbreviation can be found in the main text. For MLLMs that only support single-image input, we simply concatenate all the images vertically into one image and then input it.
@@ Appendix F Further Discussion
4B CoLVA-InternVL2-4B (Ours) 49.80 43.21 41.38 45.79 77.92 44.43 53.60 44.89 53.24
@@
Figure 9: More PCA visualizations of learned object embeddings by InternVL2-4B and our CoLVA-4B. The object embeddings are obtained by applying average pooling to the visual tokens using mask annotations. The red star represents the query object in the first image. The red dot represents the matched target in the second image. The blues dots represent other candidates.